# algorithmic

### v4

@@ 4.2 Privacy Metrics
4.3 PAC Gaussian Noise Addition Calculation
When the mutual information $MI(X;M(X))$ is insufficient to ensure PAC privacy, additional Gaussian noise $B\sim\mathcal{N}(0,\Sigma_{B})$ can be introduced to yield a reduced mutual information $MI(X;M(X)+B)$ (Xiao & Devadas, 2023). PAC privacy provides a $1-\gamma$ confidence in the noise determination for the deterministic mechanism $M$ shown in Algorithm 3, to compute the Gaussian noise covariance $\Sigma_{B}$, ensuring that $MI(X;M(X)+B)\leq\nu+\beta$. In practical terms, this means we can evaluate the privacy extent of the model by comparing the mean norm $\mathbb{E}||B||_{2}$ under the same confidence level. A smaller $\mathbb{E}||B||_{2}$ implies a smaller covariance matrix and lesser noise addition, signifying a higher privacy extent of the model.
1:  Train $M$ model with data $X$.
2:  for $k=1,2,\ldots,m$ do
3:     sample images $y^{(k)}=M(X^{(k)})$.
4:  end for
5:  Calculate empirical mean $\hat{\mu}=\frac{1}{m}\sum_{k=1}^{m}y^{(k)}$ and the empirical covariance estimation $\hat{\Sigma}=\frac{1}{m}\sum_{k=1}^{m}(y^{(k)}-\hat{\mu})(y^{(k)}-\hat{\mu})^{T}$.
6:  Apply singular value decomposition (SVD) on $\hat{\Sigma}$ and obtain the decomposition as $\hat{\Sigma}=\hat{U}\hat{\Lambda}\hat{U}^{T}$, where $\hat{\Lambda}$ is the diagonal matrix of eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{d}$.
7:  Determine the maximal index $j_{0}=\arg\max_{j}\lambda_{j}$ for those $\lambda_{j}>c$.
8:  if $\min_{1\leq j\leq j_{0},1\leq d}(\lambda_{j}-\hat{\lambda}_{j})>r\sqrt{d/c+2c}$ then then
9:     for $j=1,2,\ldots,d$ do
10:        Determine the $j$-th element of a diagonal matrix $A_{B}$ as
11:     end for
12:     Determine the Gaussian noise covariance as $\Sigma_{B}=\hat{U}A_{B}\hat{U}^{T}$.
13:  else
14:     Determine the Gaussian noise covariance as $\Sigma_{B}=\left(\sum_{j=1}^{d}\lambda_{j}+dc/(2\nu)\right)\cdot I_{d}$.
15:  end if
16:  Return Gaussian covariance matrix $\Sigma_{B}$.
@@ 5.2 Baselines
In our study, we consider DPGEN (Chen et al., 2022) and DPDM (Dockhorn et al., 2022) as baseline methods. Both of these approaches excel in synthesizing images under differential privacy (DP) constraints, and they stand out for their exceptional sample quality in comparison to other DP generative models. These models serve as important benchmarks against which we evaluate the performance and efficacy of our proposed method.
@@ 5.3 Evaluation Metrics
In our evaluation process, we validate the capability of our PAC Privacy Preserving Diffusion Model to generate high-resolution images using two key metrics: (1) the Frechet Inception Distance (Heusel et al., 2017) and (2) the Inception Score (Salimans et al., 2016). These metrics are widely recognized and utilized in the field of generative models to assess the visual fidelity of the images they produce.
Additionally, to demonstrate our model’s effectiveness in preventing privacy leakage, we employ our unique privacy metrics as outlined in Algorithm 2. We compare the mean norm of the Gaussian noise $\mathbb{E}||B||_{2}$ as detailed in Sec. 4.3. For our experiments, we have chosen the hyperparameters $\nu=\beta=0.5$ and $\gamma=0.01$. This approach allows us to comprehensively assess not just the quality of the images generated, but also the strength of privacy protection our model offers.
Table 1: Perceptual and privacy score comparisons on CelebA with image resolution $64\times 64$. In our model, we train with data on a specific label respectively, and Model-Gender means our model is trained with CelebA-Gender dataset. The same applies to the following.
P3DM-Gender P3DM-Smile DPGEN DPDM
$\varepsilon$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$
1 175$\pm$2.5 0.6$\pm$0.05 170$\pm$1.85 0.63$\pm$0.08
5 125.8$\pm$3.6 0.55$\pm$0.082 107.85$\pm$4.7 0.6$\pm$0.075 155$\pm$1.5 0.46 $\pm$ 0.076 170$\pm$2 0.57$\pm$0.06
10 44.82$\pm$4.48 0.5$\pm$0.05 37.96$\pm$1.85 0.56$\pm$0.045 39.16$\pm$0.68 0.4$\pm$0.03 117$\pm$1.5 0.47$\pm$0.04
15 40$\pm$0.5 0.45$\pm$0.05 37.9$\pm$0.6 0.5$\pm$0.025 38.5$\pm$0.73 0.4$\pm$0.06 113$\pm$1.2 0.35$\pm$0.05
$\infty$ 34.64$\pm$2.18 0.4$\pm$0.04 36.03$\pm$1.32 0.48$\pm$0.055 34.4$\pm$1.8 0.376$\pm$0.03 110$\pm$1.2 0.3$\pm$0.06
@@ 5.4 Empirical Results And Analysis
It is important to note that the $\varepsilon$ presented in our model within the tables below is merely a hyperparameter derived from the Randomized Response (RR) as indicated in Eqn. 4. This $\varepsilon$ should not be confused with the $\varepsilon$ from $\varepsilon$-Differential Privacy (DP), as our model uses PAC privacy rather than $\varepsilon$-DP. We will later illustrate how our model assures privacy through the automatic control of mutual information for a PAC privacy guarantee in this section.
@@
(c) P3DM(Ours) FID = 37.96, Privacy Score = 0.56
@@
Table 1 and Fig.2 details the evaluation results of image visual quality and privacy score on the CelebA dataset with a resolution of $64\times 64$, where each datapoint in the figure, or each entry in the table, consists of mean and standard deviation from 3 experimental results with the same epsilon and different random seeds. By examining the table, we can see that our model achieves image quality comparable to the state-of-the-art model DPGEN (Chen et al., 2022), a conclusion also supported by Fig.1.
@@
Figure 2: Privacy score and FID curve of all datapoints from different models. Each datapoint in the figure consists of mean and standard deviation from 3 experimental results with the same epsilon and different random seeds. Top-right corner is preferred. The curve from our method, pushes the frontier to the upper-right over DPGEN (Chen et al., 2022) and DPDM (Dockhorn et al., 2022).
@@
Table 2: $\mathbb{E}||B||_{2}$ to ensure $MI(X;M(X)+B)\leq 1$ on CelebA dataset
Methods $\varepsilon$ $\mathbb{E}||B||_{2}\downarrow$
P3DM-Gender $5$ 283.03$\pm$2.25
P3DM-Smile $5$ 280.60$\pm$2.57
DPGEN $5$ 281.3$\pm$2.78
P3DM-Gender $10$ 328.80$\pm$1.24
P3DM-Smile $10$ 327.98$\pm$1.45
DPGEN $10$ 329.48$\pm$1.6
DPDM $10$ 335.83$\pm$1.21
P3DM-Gender $\infty$ 332.87$\pm$1.15
P3DM-Smile $\infty$ 331.67$\pm$1.3
DPGEN $\infty$ 332.02$\pm$1.32
DPDM $\infty$ 335.83$\pm$1.56
@@
6 Limitations
The dataset attributes analyzed in our study predominantly consist of features with balanced positive and negative samples. Attributes characterized by imbalanced distributions remain underexplored. Furthermore, it should be noted that not all datasets include annotated attributes. Therefore, utilizing image-to-text models such as CLIP (Radford et al., 2021) may prove beneficial for the generation of new attribute labels. Lastly, our model faces challenges in synthesizing higher-resolution privacy-preserving images (e.g., 256x256 pixels) and handling complex datasets such as the CUB dataset (Wah et al., 2011)).

### v5

@@ 4.2 Privacy Metrics
4.3 PAC Privacy Proof of Our Model
Since the randomized response method adds Gaussian noise $Z$ in the random response mechanism, it can be proven to be PAC private using Theorem 3.1 and Theorem 3.2 in the paper:
  • Algorithm 2 can be combined and written as $H(X)+\bm{B}$, where $H(X)$ represents the RR mechanism and $\bm{B}$ denotes Gaussian noise with $\bm{B}\sim\mathcal{N}(0,\Sigma_{B})$.
  • Applying Theorem 3.2 and Theorem 3.3, we can yield a satisfied upper bound (with arbitary high confidence) for the whole process $MI(X;\mathcal{M}(X))\leq\frac{1}{2}\cdot\log\text{det}\big{(}\bm{I}_{d}+\Sigma_{\mathcal{M}(X)}\cdot\Sigma^{-1}_{\bm{B}}\big{)}$, where $\mathcal{M}(X)=H(X)+\bm{B}$.
  • Hence, $MI(X;\mathcal{M}(X))$ can be used to produce an upper bound for the attack success probability via Theorem 3.1.
What’s more, Algorithm 3 is a heuristic method that further enhances privacy. We calculate noise $B$ to measure its privacy strength, achieving the best performance as shown in Table 1. Of independent interest to ensure the PAC privacy of Algorithm 3 itself, we simply need to add noise $B$ to the final $x_{T}$ in Algorithm 3.
@@ 5.2 Baselines
In our study, we consider DPGEN (Chen et al., 2022), DPDM (Dockhorn et al., 2022) and DP-MEPF (Harder et al., 2023) as baseline methods. These approaches, except for DPGEN, which is not a rigorous DP, excel in synthesizing images under differential privacy (DP) constraints, and they stand out for their exceptional sample quality in comparison to other DP generative models. These models serve as important benchmarks against which we evaluate the performance and efficacy of our proposed method.
@@ 5.3 Evaluation Metrics
In our evaluation process, we validate the capability of our PAC Privacy Preserving Diffusion Model to generate high-resolution images using the key metric: the Frechet Inception Distance (Heusel et al., 2017). This metric is widely recognized and utilized in the field of generative models to assess the visual fidelity of the images they produce.
Additionally, to demonstrate our model’s effectiveness in preventing privacy leakage, we employ our unique privacy metrics as outlined in Algorithm 4. We compare the mean norm of the Gaussian noise $\mathbb{E}||B||_{2}$ as detailed in Theorem 3.2. For our experiments, we have chosen the hyperparameters $\nu=\beta=0.5$ and $\gamma=0.01$. This approach allows us to comprehensively assess not just the quality of the images generated, but also the strength of privacy protection our model offers. It is important to note that hyperparameter tuning does incur an additional privacy cost, albeit modest (Liu & Talwar, 2019; Papernot & Steinke, 2022; Ding & Wu, 2023; Xiang et al., 2024). For simplicity we ignore the privacy overhead due to hyperparameter tuning for all methods.
@@ 5.4 Empirical Results And Analysis
It is important to note that the $\xi$ and $\tau$ presented within the tables below is merely a hyperparameter derived from the Randomized Response (RR) as indicated in Eqn. 12. This $\xi$ and $\tau$ should not be confused with the $\varepsilon$ from $\varepsilon$-Differential Privacy (DP), where $\xi$ indicates the model is PAC private, and $\tau$ indicates the model is neither PAC private nor DP. We will later illustrate how our model assures privacy through the automatic control of mutual information for a PAC privacy guarantee in this section.
@@
(c) DP-MEPF FID = 57.5, Privacy Score = 0.4
(d) P3DM (Ours) FID = 37.96, Privacy Score = 0.56
@@
Fig. 3 details the evaluation results of image visual quality and privacy score on the CelebA dataset with a resolution of $64\times 64$, where each datapoint in the figure, or each entry in the table, consists of mean and standard deviation from 3 experimental results with the same epsilon and different random seeds. By examining the table, we can see that our model achieves image quality comparable to the state-of-the-art model DPGEN (Chen et al., 2022), a conclusion also supported by Fig.1.
@@
Figure 3: Privacy score and FID curve of all datapoints from different models. Each datapoint in the figure consists of mean and standard deviation from 3 experimental results with the same epsilon and different random seeds. Top-right corner is preferred. The curve from our method, pushes the frontier to the upper-right over DPGEN (Chen et al., 2022), DPDM (Dockhorn et al., 2022) and DP-MEPF (Harder et al., 2023). For more details related to data points in this figure, please refer to Appendix C.
Table 1: Estimated level of noise $B$ to approximately ensure PAC privacy
Methods DP PAC Privacy Heuristic $\mathbb{E}||B||_{2}\downarrow$
P3DM-Gender $\xi=5$ 283.03$\pm$2.25
P3DM-Smile $\xi=5$ 280.60$\pm$2.57
DPGEN $\tau=5$ 281.3$\pm$2.78
DP-MEPF $\varepsilon=5$ 325.6$\pm$3.62
P3DM-Gender $\xi=10$ 328.80$\pm$1.24
P3DM-Smile $\xi=10$ 327.98$\pm$1.45
DPGEN $\tau=10$ 329.48$\pm$1.6
DPDM $\varepsilon=10$ 335.83$\pm$1.21
DP-MEPF $\varepsilon=10$ 330.75$\pm$1.33
P3DM-Gender $\xi=\infty$ 332.87$\pm$1.15
P3DM-Smile $\xi=\infty$ 331.67$\pm$1.3
DPGEN $\tau=\infty$ 332.02$\pm$1.32
DPDM $\varepsilon=\infty$ 335.83$\pm$1.56
DP-MEPF $\varepsilon=\infty$ 333.48$\pm$1.79
@@ 6 Conclusion
Appendix A Proof of Theorem 3.1 (Xiao & Devadas, 2023)
Proof.
To start, we need the following lemmas.
Lemma A.1.
Given any $f$-divergence $\mathcal{D}_{f}(\cdot\|\cdot)$, and three Bernoulli distributions $\bm{1}_{a}$, $\bm{1}_{b}$ and $\bm{1}_{c}$ of parameters $a$, $b$ and $c$, respectively, where $0\leq a\leq b\leq c\leq 1$. Then, $\mathcal{D}_{f}(\bm{1}_{a}\|\bm{1}_{b})\leq\mathcal{D}_{f}(\bm{1}_{a}\|\bm{1}_{c}).$
Proof.
By the definition, $g(x)=\mathcal{D}_{f}(\bm{1}_{a}\|\bm{1}_{x})=xf(\frac{a}{x})+(1-x)f(\frac{1-a}{1-x})$ and we want to show $g(x)$ is non-decreasing for $x\geq a$. With some calculation, $g^{\prime}(x)=\big{(}f(\frac{a}{x})-\frac{a}{x}f^{\prime}(\frac{a}{x})\big{)}-\big{(}f(\frac{1-a}{1-x})-\frac{1-a}{1-x}f^{\prime}(\frac{1-a}{1-x})\big{)}.$ It is noted that $\frac{a}{x}\leq\frac{1-a}{1-x}$ for $x\geq a$. Thus, to show $g^{\prime}(x)\geq 0$ for $x\geq a$, it suffices to show $t(y)=f(y)-yf^{\prime}(y)$ is non-increasing with respect to $y\in[0,1]$. On the other hand, $t^{\prime}(y)=f^{\prime}(y)-f^{\prime}(y)-yf^{\prime\prime}(y)\leq 0$ due to the convex assumption of $f$. Therefore, the claim holds. ∎∎
Lemma A.2 (Data Processing Inequality Sason & Verdú (2016)).
Consider a channel that produces $Z$ given $Y$ based on the law described as a conditional distribution $\mathsf{P}_{Z|Y}$. If $\mathsf{P}_{Z}$ is the distribution of $Z$ when $Y$ is generated by $\mathsf{P}_{Y}$, and $\mathsf{Q}_{Z}$ is the distribution of $Z$ when $Y$ is generated by $\mathsf{Q}_{Y}$, then for any f-divergence $\mathcal{D}_{f}$,
$\mathcal{D}_{f}(\mathsf{P}_{Z}\|\mathsf{Q}_{Z})\leq\mathcal{D}_{f}(\mathsf{P}_{Y}\|\mathsf{Q}_{Y}).$
Now, we return to prove Theorem 3.1. First, we have the observation that for a random variable $X^{\prime}\in\mathcal{X}^{*}$ in an arbitrary distribution but independent of $X$, $\delta^{\rho}_{o}\leq\Pr_{X^{\prime}\perp X}\big{(}\rho(X^{\prime},X)=1\big{)},$ since $\delta^{\rho}_{o}$ is the minimum failure probability achieved by optimal a priori estimation. Here, $a\perp b$ represents that $a$ is independent of $b$. Let the indicator be a function that for two random variables $a$ and $b$, $\bm{1}(a,b)=1$ if $\rho(a,b)=1$, otherwise 00. Apply Lemma A.1 and Lemma A.2, where we view $\bm{1}(\cdot,\cdot)$ as a post-processing on $(X,\tilde{X})$ and $(X,{X}^{\prime})$, respectively, we have that
$\mathcal{D}_{f}\big{(}\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}}\big{)}\leq\mathcal{D}_{f}\big{(}\bm{1}(X,\tilde{X})\|\bm{1}(X,X^{\prime})\big{)}\leq\mathcal{D}_{f}\big{(}\mathsf{P}_{X,\tilde{X}}\|\mathsf{P}_{X,X^{\prime}}\big{)}=\mathcal{D}_{f}\big{(}\mathsf{P}_{X,\tilde{X}}\|\mathsf{P}_{X}\otimes\mathsf{P}_{X^{\prime}}\big{)}.$
On the other hand, we know $X\to\mathcal{M}(X)\to\tilde{X}$ forms a Markov chain, where the adversary’s estimation $\tilde{X}$ is dependent on observation $\mathcal{M}(X)$. Let the adversary’s strategy be some operator $g_{adv}$ where $\tilde{X}=g_{adv}(\mathcal{M}(X))$. Therefore, we can apply the data processing inequality again, where
$\mathcal{D}_{f}\big{(}\mathsf{P}_{X,\tilde{X}}\|\mathsf{P}_{X,X^{\prime}}\big{)}\leq\mathcal{D}_{f}\big{(}\mathsf{P}_{X,\mathcal{M}(X)}\|\mathsf{P}_{X,W}\big{)}=\mathcal{D}_{f}\big{(}\mathsf{P}_{X,\mathcal{M}(X)}\|\mathsf{P}_{X}\otimes\mathsf{P}_{W}\big{)}.$
Here, $X^{\prime}=g_{adv}(W)$ and $W$ is still independent of $X$. Since the above inequalities hold for arbitrarily distributed $X^{\prime}$ once it is independent of $X$, $W$ could also be an arbitrary random variable on the same support domain as $\mathcal{M}(X)$ and independent of $X$. Therefore,
$\mathcal{D}_{f}\big{(}\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}}\big{)}\leq\inf_{\mathsf{P}_{W}}\mathcal{D}_{f}\big{(}\mathsf{P}_{X,\mathcal{M}(X)}\|\mathsf{P}_{X}\otimes\mathsf{P}_{W}\big{)}=\inf_{\mathsf{P}_{W}}\mathcal{D}_{f}\big{(}\mathsf{P}_{\mathcal{M}(X)|X}\|\mathsf{P}_{W}|\mathsf{P}_{X}).$
Here, we use $|\mathsf{P}_{X}$ to denote that it is conditional on $X$ in a distribution $\mathsf{P}_{X}$. In particular, if we select $\mathsf{P}_{W}$ to be the distribution of $\mathcal{M}(X)$, and take $\mathsf{D}_{f}$ to be KL-divergence, we have $\mathcal{D}_{KL}\big{(}\bm{1}_{\delta}\|\bm{1}_{\delta^{\rho}_{o}}\big{)}\leq\mathsf{MI}(X;\mathcal{M}(X)).$
Appendix B Proof of Theorem 3.2 (Xiao & Devadas, 2023)
Proof.
For $\mathsf{MI}(X;\mathcal{M}(X)+B)$, we have
$\mathsf{MI}(X;\mathcal{M}(X)+B)$ (17)
[…]
Given the definition of mutual information, we first apply the results of Gaussian approximation Pinsker et al. (1995), where $Gau(A)$ represents a (multivariate) Gaussian variable with the same mean and (co)variance as those of $A$. Then, we drop a negative term to obtain the final inequality of (17). Next, focusing on the integral (first) term in (17),
$\mathcal{D}_{KL}(\mathsf{P}_{\mathcal{M}(X_{0})+B}\|\mathsf{P}_{B})=\frac{1}{2}\cdot(\mathcal{M}(X_{0}))^{T}\Sigma^{-1}_{\bm{B}}(\mathcal{M}(X_{0})),$
and therefore
[…]
As for the second term in the last equation of (17), we have the covariance of $\mathcal{M}(X)$ equals $\Sigma_{\mathcal{M}(X)}=\mathbb{E}_{X}\big{[}(\mathcal{M}(X)-\mathbb{E}[\mathcal{M}(X)])(\mathcal{M}(X)-\mathbb{E}[\mathcal{M}(X)])^{T}\big{]}$, while the mean $\mu_{\mathcal{M}(X)}=\mathbb{E}[\mathcal{M}(X)]$. The KL divergence between two multivariate Gaussians has a closed form, where $\mathcal{D}_{KL}(\mathsf{P}_{Gau(\mathcal{M}(X)+B)}\|\mathsf{P}_{B})$ equals
$\mathcal{D}_{KL}$ $(\mathsf{P}_{Gau(\mathcal{M}(X)+B)}\|\mathsf{P}_{B})=\frac{1}{2}\cdot\big{(}\text{Trace}(\Sigma_{\mathcal{M}(X)}\cdot\Sigma^{-1}_{\bm{B}})$ (18)
$+\mathbb{E}_{X}[\mathcal{M}(X)]^{T}\Sigma^{-1}_{\bm{B}}\mathbb{E}_{X}[\mathcal{M}(X)]-\log\text{det}(\bm{I}_{d}+\Sigma_{\mathcal{M}(X)}\Sigma^{-1}_{\bm{B}})\big{)}.$
On the other hand, note that
$\mathbb{E}_{X}\big{[}(\mathcal{M}(X))^{T}\Sigma^{-1}_{\bm{B}}(\mathcal{M}(X))\big{]}-\mathbb{E}_{X}\big{[}\mathcal{M}(X)\big{]}^{T}\Sigma^{-1}_{\bm{B}}\mathbb{E}_{X}\big{[}\mathcal{M}(X)\big{]}-\text{Trace}(\Sigma_{\mathcal{M}(X)}\cdot{\Sigma}^{-1}_{\bm{B}})$ (19)
$=\text{Trace}\big{(}\mathbb{E}[\mathcal{M}(X)-\mathbb{E}[\mathcal{M}(X)]]\cdot\Sigma^{-1}_{\bm{B}}\cdot\mathbb{E}[\mathcal{M}(X)-\mathbb{E}[\mathcal{M}(X)]]^{T}\big{)}-\text{Trace}\big{(}\Sigma_{\mathcal{M}(X)}\Sigma^{-1}_{\bm{B}}\big{)}$
$=\text{Trace}\big{(}\mathbb{E}[\mathcal{M}(X)-\mathbb{E}[\mathcal{M}(X)]]\cdot\mathbb{E}[\mathcal{M}(X)-\mathbb{E}[\mathcal{M}(X)]]^{T}\cdot\Sigma^{-1}_{\bm{B}}-\Sigma_{\mathcal{M}(X)}\Sigma^{-1}_{\bm{B}}\big{)}$
$=\text{Trace}\big{(}\Sigma_{\mathcal{M}(X)}\cdot\Sigma^{-1}_{\bm{B}}-\Sigma_{\mathcal{M}(X)}\Sigma^{-1}_{\bm{B}}\big{)}=0.$
In (19), we use the following facts that for two arbitrary vectors $v_{1},v_{2}\in\mathbb{R}^{d}$, $(v_{1})^{T}v_{2}=\text{Trace}(v_{1}\cdot(v_{2})^{T})$, and for two arbitrary matrices $A_{1},A_{2}\in\mathbb{R}^{d\times d}$, $\text{Trace}(A_{1}A_{2})=\text{Trace}(A_{2}A_{1})$. Therefore, putting it all together, we have a simplified form of the right hand of (17), where
$\mathsf{MI}(X;\mathcal{M}(X)+B)\leq\frac{\log\text{det}(\bm{I}_{d}+\Sigma_{\mathcal{M}(X)}\cdot\Sigma^{-1}_{\bm{B}})}{2}.$ (20)
Appendix C Detailed Data in Figure 3
Table 2: Perceptual and privacy score comparisons on CelebA with image resolution $64\times 64$. In our model, we train with data on a specific label respectively, and Model-Gender means our model is trained with CelebA-Gender dataset. $\xi$ indicates a hyperparameter derived from the Randomized Response (RR).
DPGEN P3DM-Gender P3DM-Smile
$\xi$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$
1 175$\pm$2.5 0.6$\pm$0.05 170$\pm$1.85 0.63$\pm$0.08
5 155$\pm$1.5 0.46$\pm$0.076 125.8$\pm$3.6 0.55$\pm$0.082 107.85$\pm$4.7 0.6$\pm$0.075
10 39.16$\pm$0.68 0.4$\pm$0.03 44.82$\pm$4.48 0.5$\pm$0.05 37.96$\pm$1.85 0.56$\pm$0.045
15 38.5$\pm$0.73 0.4$\pm$0.06 40$\pm$0.5 0.45$\pm$0.05 37.9$\pm$0.6 0.5$\pm$0.025
$\infty$ 34.4$\pm$1.8 0.376$\pm$0.03 34.64$\pm$2.18 0.4$\pm$0.04 36.03$\pm$1.32
Table 3: Perceptual and privacy score comparisons on CelebA with image resolution $64\times 64$. In our model, we train with data on a specific label respectively, and Model-Gender means our model is trained with CelebA-Gender dataset. $\varepsilon$ represents the DP parameter of baselines.
DPDM DP-MEPF
$\varepsilon$ FID $\downarrow$ Privacy Score $\uparrow$ FID $\downarrow$ Privacy Score $\uparrow$
1 67.5$\pm$1.45 0.48$\pm$0.02
5 170$\pm$2 0.57$\pm$0.06 61.2$\pm$1.8 0.45$\pm$0.03
10 117$\pm$1.5 0.47$\pm$0.04 57.5$\pm$1.2 0.4$\pm$0.04
15 113$\pm$1.2 0.35$\pm$0.05 55.8$\pm$1.4 0.38$\pm$0.02
$\infty$ 110$\pm$1.2 0.3$\pm$0.06 52$\pm$2.3 0.34$\pm$0.025